I need help making a system of equations based off of a word problem. Here is the problem: Mary and John save their quarters. Mary starts out with 34 quarters and saves 8 quarters a month. John starts out with 2 quarters and saves 16 quarters a month. In how many months will they have the same number of quarters?
Answer by Alan3354(69443)

Mary and John save their quarters. Mary starts out with 34 quarters and saves 8 quarters a month. John starts out with 2 quarters and saves 16 quarters a month. In how many months will they have the same number of quarters?
---------------
M = 34 + 8t (t = months)
J = 2 + 16t
When they're equal,
34+8t = 2+16t
32 = 8t
t = 4 months